I need help on finding the zeros to this function.... f(x)=x^2+6x+18
Answer by rothauserc(4718)

use the quadratic formula ((-b+ or -sqrt(b^2-4ac))/2a) to solve
x = (-6+sqrt(36-4*18)) / 2 = (-6+6i) / 2 = 3i-3 = 3(i-1) where i = sqrt(-1)
x = (-6-sqrt(36-4*18)) / 2 = (-6-6i) / 2 = -3i-3 = -3(i+1)